### v1

@@ 4.4 Predicting relaxation quantities
Assuming PaEGNN comprises $T$ layers, we define $\bm{h}_{ij}=\gamma\big{(}\lambda(\|\vec{\bm{r}}_{ji}\|)\oplus\bm{p}(k_{1},k_{2},k_{3})\big{)}$. The prediction of a pair-wise distance $\hat{d}_{ij}$ and its corresponding local uncertainty $\hat{u}_{ij}$ for the edge $e_{ij,(k_{1},k_{2},k_{3})}$ is formulated as:
$\hat{d}_{ij}=|f_{d}(\bm{W}_{n}\bm{x}_{i}^{(T)}\oplus\bm{W}_{n}\bm{x}_{i}^{(T)}\oplus\bm{h}_{ij})|$ (8)
$\hat{u}_{ij}=f_{u}(\bm{W}_{n}\bm{x}_{i}^{(T)}\oplus\bm{W}_{n}\bm{x}_{i}^{(T)}\oplus\bm{h}_{ij})$ (9)
where $\bm{W}_{n}\in\mathbb{R}^{F\times F}$ is a learnable matrix, and $f_{d},f_{u}:\mathbb{R}^{3F}\rightarrow\mathbb{R}$ are linear maps. Using Eqn. (8) and (9), we can predict both the interatomic distances in the relaxed structure and the displacements between the initial and relaxed structures, along with their respective local uncertainties.
DeepRelax predicts the lattice matrix as follows:
@@ 4.5 Uncertainty-aware loss function
To quantify aleatoric uncertaintykendall2017uncertainties ; yu2022uncertainty , DeepRelax is trained to minimize an uncertainty-aware loss $\mathcal{L}$, comprising interatomic distance loss $\mathcal{L}_{i}$, displacement loss $\mathcal{L}_{d}$, and lattice loss $\mathcal{L}_{l}$:
$\mathcal{L}_{i}=\sum_{e_{ij,(k_{1},k_{2},k_{3})}\in\mathcal{E}}\log(2\hat{u}_{ij})+\frac{|\hat{d}_{ij}-d_{ij}|}{\hat{u}_{ij}}$ (11)
$\mathcal{L}_{d}=\sum_{e_{ij,(0,0,0)}\in\mathcal{E}}\log(2\hat{u}_{ij})+\frac{|\hat{d}_{ij}-d_{ij}|}{\hat{u}_{ij}}$ (12)
$\mathcal{L}_{l}=|\hat{\bm{L}}-\bm{L}|$ (13)
$\mathcal{L}=\mathcal{L}_{i}+\mathcal{L}_{d}+\mathcal{L}_{l}$ (14)
In these expressions, $d_{ij}$ and $\bm{L}$ represent the ground truth distance and lattice matrix of the relaxed structure, respectively. The edges $e_{ij,(k_{1},k_{2},k_{3})}\in\mathcal{E}$ pertain to interatomic distance predictions, whereas $e_{ij,(0,0,0)}\in\mathcal{E}$ denotes edges used for displacement predictions within the unit cell, discounting PBCs.
Minimizing $\mathcal{L}_{i}$ and $\mathcal{L}_{d}$ is mathematically equal to maximize the Laplace log-likelihood as:
$\log p(x|\mu,b)=-\log(2b)-\frac{|x-\mu|}{b}$ (15)
where $\mu$ and $b$ are the mean and variance of the Laplace log-likelihood, respectively. Therefore, minimizing $\mathcal{L}_{i}$ and $\mathcal{L}_{d}$ encourages the output distances to follow a distribution akin to the Laplace log-likelihood, where $\hat{d}_{ij}$ is the mean and $\hat{u}_{ij}$ is the variance. Consequently, a larger $\hat{u}_{ij}$ indicates greater local uncertainty in the prediction, and vice versa.
@@ 4.6 Numerical Euclidean distance geometry solver
For simplicity, $\bm{\hat{R}}(u)$ and $\bm{\hat{R}}(v)$ are typically written as $\bm{\hat{R}}_{u}$ and $\bm{\hat{R}}_{v}$.
@@
In our context, we aim to find a coordinate matrix $\bm{\hat{R}}\in\mathbb{R}^{N\times 3}$ for a system of $N$ atoms in three-dimensional space, meeting the constraints imposed by $\hat{d}_{ij}$, $\hat{u}_{ij}$, and $\hat{\bm{L}}$. Specifically, we first define an upper bound and lower bound using $\hat{d}_{ij}$, $\hat{u}_{ij}$ as follows:
$\hat{d}_{ij}^{u}=\hat{d}_{ij}+\exp(\hat{u}_{ij})$ (18)
$\hat{d}_{ij}^{l}=\hat{d}_{ij}-\exp(\hat{u}_{ij})$ (19)
@@
For each edge $e_{ij,(k_{1},k_{2},k_{3})}$, the location of node $j$ is dictated by $k_{1}\hat{\bm{l}}_{1}+k_{2}\hat{\bm{l}}_{2}+k_{3}\hat{\bm{l}}_{3}$, where $\hat{\bm{l}}_{1}$, $\hat{\bm{l}}_{2}$, $\hat{\bm{l}}_{3}$ are predicted lattice vectors. The BED loss only penalizes coordinate pairs whose distances fall outside the lower and upper bounds, thus mitigating the impact of less accurate predictions. In our work, we use Adam optimizerkingma2014adam to minimize $\mathcal{L}_{g}$.
@@ 4.7 Uncertainty quantification
We use global uncertainty to assess the reliability of the predicted structure. This involves training an ensemble of $T$ independent model replicates, with $M=5$ used in this study, a method that has been extensively validated for effective uncertainty estimation yu2022uncertainty ; luo2023calibrated . The $T$ model replicates had the same neural network architectures and hyperparameters, but the learnable parameters were initialized with different random seeds. For the $t$-th model replicate, let $\hat{d}_{ij}(t)$ denote the predicted distance and $\hat{u}_{ij}(t)$ the local uncertainty. The global uncertainty, $\hat{U}$, is computed as follows:
$\hat{U}_{a}=\bigg{(}\frac{1}{N}\sum_{\begin{subarray}{c}e_{ij,(k_{1},k_{2},k_{3})}\in\mathcal{E}\\ e_{ij,(0,0,0)}\in\mathcal{E}\end{subarray}}\Big{(}\frac{1}{T}\sum_{t=1}^{T}\exp\big{(}\hat{u}_{ij}(t)\big{)}\Big{)}\bigg{)}^{2}$ (21)
@@
$\hat{U}_{e}=\frac{1}{N}\sum_{\begin{subarray}{c}e_{ij,(k_{1},k_{2},k_{3})}\in\mathcal{E}\\ e_{ij,(0,0,0)}\in\mathcal{E}\end{subarray}}\Big{(}\frac{1}{T}\sum_{t=1}^{T}\big{(}\hat{d}_{ij}(t)-\bar{d}_{ij}\big{)}^{2}\Big{)}$ (23)
$\hat{U}=\hat{U}_{a}+\hat{U}_{e}$ (24)
In these equations, $N$ represents the total number of distance pairs. The terms $\hat{U}_{a}$ and $\hat{U}_{e}$ are known as aleatoric and epistemic uncertainties, respectivelykendall2017uncertainties .
@@ 4.8 Implementation details
The DeepRelax model is implemented using PyTorch. Experiments are conducted on an NVIDIA RTX A6000 with 48 GB of memory. The training objective is to minimize Eqn. (14). We use the AdamW optimizerloshchilov2017decoupled with a learning rate of 0.0001 to update the model’s parameters. Additionally, we implement a learning rate decay strategy, reducing the learning rate if there is no improvement in a specified metric for a duration of 5 epochs.
@@ 4.9 DFT calculations
In our study, DFT calculations are performed using the Vienna Ab initio Simulation Package (VASP)kresse1996efficient , employing the generalized gradient approximation (GGA) with the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functionalperdew1996generalized . We set the energy cut-off to 550 eV, the energy convergence criterion to $1.0\times 10^{-5}$ eV, and the K-point mesh to a $9\times 9\times 9$ grid, ensuring precise total energy calculations. The effective on-site Coulomb interactions (U value) of Mn 3d orbital is chosen as 3.9 eV, aligning with that used in Cryslatorkim2023structure . For high-throughput self-consistent calculations, we utilize the AiiDA computational frameworkhuber2020aiida ; uhrin2021workflows .
@@ 4.10.1 MAE of coordinate
The MAE of coordinates assesses the structural difference between the predicted and actual relaxed structures. It is defined as:
$\Delta_{\rm coord}=\sum_{v_{i}\in\mathcal{G}}\lvert\hat{\bm{r}}_{i}-\vec{\bm{r}}_{i}\rvert$ (25)
where $\hat{\bm{r}}_{i}$, $\vec{\bm{r}}_{i}$ represent the predicted and ground truth Cartesian coordinates, respectively.
@@ 4.10.2 MAE of bond length
$\Delta_{\rm bond}=\sum_{e_{ij,(k_{1},k_{2},k_{3})}\in\mathcal{E}}|\hat{d}_{ij}-d_{ij}|$ (26)
where $\hat{d}_{ij}$, $d_{ij}$ are the predicted and ground interatomic distances.
@@ 4.10.3 MAE of lattice
This metric calculates the error in predicting the lattice parameters:
$\Delta_{\rm lattice}=|\hat{\bm{L}}-\bm{L}|$ (27)
where $\hat{\bm{L}}$ and $\bm{L}$ are the predicted and ground lattice parameters.
@@ 4.10.4 MAE of cell volume
$\Delta_{\rm volume}=|\hat{\bm{l}}_{1}\cdot(\hat{\bm{l}}_{2}\times\hat{\bm{l}}_{3})|-|\vec{\bm{l}}_{1}\cdot(\vec{\bm{l}}_{2}\times\vec{\bm{l}}_{3})|$ (28)
@@ 4.10.5 Match rate
We utilize the ’Structure_matcher’ function from the Pymatgen packageong2013python to compare the predicted relaxed structure with the ground truth relaxed structure. Default parameters are used for this function (ltol=0.2, stol=0.3) to ensure consistent and objective comparisons.
Conflicts of interest
The authors declare that they have no competing interests.
Acknowledgements
We acknowledge Prof. Kristian Sommer Thygesen and Peder Lyngby for their generous provision of the C2DB database, complete with both initial and final structures. This work was supported by the National Natural Science Foundation of China (Grant No. 62176272), Research and Development Program of Guangzhou Science and Technology Bureau (No. 2023B01J1016), Key-Area Research and Development Program of Guangdong Province (No. 2020B1111100001), Singapore MOE Tier 1 (No. A-8001194-00-00), and Singapore MOE Tier 2 (No. A-8001872-00-00).

### v2

@@ 4.4 Predicting relaxation quantities
Assuming PaEGNN comprises $T$ layers, and we define bond feature $\bm{h}_{ij}=\gamma\big{(}\lambda(\|\vec{\bm{r}}_{ij}\|)\oplus\bm{p}(k_{1},k_{2},k_{3})\big{)}$, where $\gamma$ is a linear projection mapping the concatenated feature back to $F$-dimensional space. The prediction of a pairwise distance $\hat{d}_{ij}$ for the edge $e_{ij,(k_{1},k_{2},k_{3})}$ is formulated as:
$\hat{d}_{ij}=|f_{d}(\bm{W}_{d}\bm{x}_{i}^{(T)}\oplus\bm{W}_{d}\bm{x}_{j}^{(T)}\oplus\bm{h}_{ij})|$ (8)
where $\bm{W}_{d}\in\mathbb{R}^{F\times F}$ is a learnable matrix, and $f_{d}:\mathbb{R}^{3F}\rightarrow\mathbb{R}$ is linear maps. Using Eqn. (8), we can predict both the interatomic distances in the relaxed structure and the displacements between the initial and relaxed structures. Additionally, DeepRelax predicts the lattice matrix of the relaxed structure as follows:
@@ 4.5 Uncertainty-aware loss function
In real scenarios, each predicted distance is subject to inherent noise (e.g., measurement errors or human labeling errors). To capture this uncertainty, we can model the pairwise distances as random variables following a Laplace distribution, i.e., $d_{ij}\sim\text{Laplace}(\hat{d}_{ij},\hat{b}_{ij})$. Here, $\hat{d}_{ij}$ and $\hat{b}_{ij}$ are the location parameter and scale parameter, respectively. In our application, $\hat{d}_{ij}$ represents the predicted distance, and $\hat{b}_{ij}$ represents the associated bond-level data uncertainty. The scale parameter $\hat{b}_{ij}$ is predicted as follows:
$\hat{b}_{ij}=f_{b}(\bm{W}_{b}\bm{x}_{i}^{(T)}\oplus\bm{W}_{b}\bm{x}_{j}^{(T)}\oplus\bm{h}_{ij})$ (10)
where $\bm{W}_{b}\in\mathbb{R}^{F\times F}$ is a learnable matrix, and $f_{b}:\mathbb{R}^{3F}\rightarrow\mathbb{R}$ is a linear map.
To train DeepRelax such that its output follows the assumed Laplace distribution, we propose an uncertainty-aware loss $\mathcal{L}_{u}$, which comprises interatomic distance loss $\mathcal{L}_{i}$ and displacement loss $\mathcal{L}_{d}$:
$\mathcal{L}_{i}=\sum_{e_{ij,(k_{1},k_{2},k_{3})}\in\mathcal{E}}\log(2\hat{b}_{ij})+\frac{|d_{ij}-\hat{d}_{ij}|}{\hat{b}_{ij}}$ (11)
$\mathcal{L}_{d}=\sum_{e_{ij,(0,0,0)}\in\mathcal{E}}\log(2\hat{b}_{ij})+\frac{|d_{ij}-\hat{d}_{ij}|}{\hat{b}_{ij}}$ (12)
$\mathcal{L}_{u}=\mathcal{L}_{i}+\mathcal{L}_{d}$ (13)
In these expressions, $d_{ij}$ represents the ground truth distance. The edges $e_{ij,(k_{1},k_{2},k_{3})}\in\mathcal{E}$ pertain to interatomic distance predictions, whereas $e_{ij,(0,0,0)}\in\mathcal{E}$ denotes edges used for displacement predictions within the unit cell, discounting PBCs. In essence, $\mathcal{L}_{i}$ and $\mathcal{L}_{d}$ represent the negative log-likelihood of the Laplace distribution, thereby capturing the data uncertainty. Consequently, a larger $\hat{b}_{ij}$ indicates greater bond-level data uncertainty in the prediction, and vice versa. The total loss $\mathcal{L}$ is consist of $\mathcal{L}_{u}$ and a lattice loss $\mathcal{L}_{l}$:
$\mathcal{L}_{l}=\sum|\hat{\bm{L}}-\bm{L}|$ (14)
$\mathcal{L}=\mathcal{L}_{u}+\mathcal{L}_{l}$ (15)
where $\bm{L}$ represents the ground lattice matrix of the relaxed structure.
@@ 4.6 Numerical Euclidean distance geometry solver
For simplicity in notation, $\bm{\hat{R}}(u)$ and $\bm{\hat{R}}(v)$ are typically written as $\bm{\hat{R}}_{u}$ and $\bm{\hat{R}}_{v}$.
@@
In our specific application, we aim to find a coordinate matrix $\bm{\hat{R}}\in\mathbb{R}^{N\times 3}$ for a system of $N$ atoms in three-dimensional space, meeting the constraints imposed by $\hat{d}_{ij}$, $\hat{b}_{ij}$, and $\hat{\bm{L}}$. Specifically, we first define an upper bound and a lower bound using $\hat{d}_{ij}$, $\hat{b}_{ij}$ as follows:
$\hat{d}_{ij}^{u}=\hat{d}_{ij}+\exp(\hat{b}_{ij})$ (18)
$\hat{d}_{ij}^{l}=\hat{d}_{ij}-\exp(\hat{b}_{ij})$ (19)
@@
For each edge $e_{ij,(k_{1},k_{2},k_{3})}$, the location of node $v_{j}$ is dictated by $k_{1}\hat{\bm{l}}_{1}+k_{2}\hat{\bm{l}}_{2}+k_{3}\hat{\bm{l}}_{3}$, where $\hat{\bm{l}}_{1}$, $\hat{\bm{l}}_{2}$, $\hat{\bm{l}}_{3}$ are predicted lattice vectors. The BED loss only penalizes coordinate pairs whose distances fall outside the lower and upper bounds, thus mitigating the impact of less accurate predictions. In our work, we use Adam optimizer to minimize $\mathcal{L}_{g}$.
@@ 4.7 Uncertainty quantification
We initially quantify bond-level uncertainties and subsequently aggregate these to determine the system-level uncertainty of the predicted structure. The bond-level uncertainty can be further decomposed into data uncertainty and model uncertainty. Data uncertainty arises from the inherent randomness in the data, while model uncertainty arises from a lack of knowledge about the best model to describe the data gawlikowski2023survey .
We employ ensemble-based uncertainty techniques yu2022uncertainty ; luo2023calibrated , which involve training an ensemble of $T$ independent model replicates, with $T=5$ used in this study. The $T$ model replicates have the same neural network architectures and hyperparameters, but the learnable parameters are initialized with different random seed. For the $t$-th model replicate, let $\hat{d}_{ij}(t)$ denote the predicted distance, $\hat{b}_{ij}(t)$ the associated bond-level data uncertainty, and $\hat{w}_{ij}(t)$ the associated bond-level model uncertainty. Model uncertainty for each pair is calculated as the deviation from the mean predicted distance $\bar{d}_{ij}$:
$\hat{w}_{ij}(t)=\lvert\hat{d}_{ij}(t)-\bar{d}_{ij}\rvert$ (21)
where the mean predicted distance $\bar{d}_{ij}$ is given by:
@@
The total bond-level uncertainty $\hat{U}_{ij}$ is the sum of the exponential of the data uncertainties and the model uncertainties across $T$ models:
$\hat{U}_{ij}=\frac{1}{T}\sum_{t=1}^{T}\Big{(}\exp(\hat{b}_{ij}(t))+\hat{w}_{ij}(t)\Big{)}$ (23)
Finally, the system-level uncertainty $\hat{U}$ is computed as the average of all bond-level uncertainties:
$\hat{U}=\frac{1}{N}\sum_{\begin{subarray}{c}e_{ij,(k_{1},k_{2},k_{3})}\in\mathcal{E}\\ e_{ij,(0,0,0)}\in\mathcal{E}\end{subarray}}\hat{U}_{ij}$ (24)
where $N$ represents the total number of evaluated pairs.
@@ 4.8 Implementation details
The DeepRelax model is implemented using PyTorch. Experiments are conducted on an NVIDIA RTX A6000 with 48 GB of memory. The training objective is to minimize Eqn. (15). We use the AdamW optimizer with a learning rate of 0.0001 to update the model’s parameters. Additionally, we implement a learning rate decay strategy, reducing the learning rate if there is no improvement in a specified metric for a duration of 5 epochs.
@@ 4.9 DFT calculations
In our study, DFT calculations are performed using the Vienna Ab initio Simulation Package (VASP) kresse1996efficient , employing the generalized gradient approximation (GGA) with the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional. All VASP calculations are performed using the electronic minimization algorithm “all band simultaneous update of orbitals” (ALGO=All), with a cut-off energy of 550 eV, an energy convergence criterion of $1.0\times 10^{-5}$ eV, and a Gaussian smearing width of 0.02 eV. For the X-Mn-O dataset, we run the self-consistent calculation to obtain the total energy without spin polarization. The K-point mesh is a $9\times 9\times 9$ grid, ensuring precise total energy calculations. The effective on-site Coulomb interactions (U value) of Mn 3d orbital is chosen as 3.9 eV, aligning with that used in Cryslator kim2023structure . For the MP dataset, the self-consistent is running with a $5\times 5\times 5$ K-point mesh for structures containing fewer than 60 atoms and $3\times 3\times 3$ for those with more than 60 atoms. Spin polarization is applied to structures exhibiting magnetism to enhance the convergence of total energy calculations. For layered vdW crystals, we performed DFT calculations with van der Waals corrections (DFT-D3 Grimme method). For $\rm MoS_{2}$ structures with defects, the structure is relaxed until the interatomic force is smaller than $0.05\ \rm{eV/\AA}$. Spin polarization is included following previous studies huang2023unveiling ; kazeev2023sparse . These high-throughput self-consistent and structural relaxation calculations are implemented utilizing the AiiDA computational framework uhrin2021workflows .
@@ 4.10.1 MAE of coordinate
The MAE of coordinates assesses the structural difference between the predicted and DFT-relaxed structures. It is defined as:
$\Delta_{\rm coord}=\frac{1}{N}\sum_{v_{i}\in\mathcal{G}}\lvert\hat{\bm{r}}_{i}-\vec{\bm{r}}_{i}\rvert$ (25)
where $N$ represents the total number of nodes in $\mathcal{G}$, $\hat{\bm{r}}_{i}$ and $\vec{\bm{r}}_{i}$ represent the predicted and ground truth Cartesian coordinates, respectively.
@@ 4.10.2 MAE of bond length
$\Delta_{\rm bond}=\frac{1}{M}\sum_{e_{ij,(k_{1},k_{2},k_{3})}\in\mathcal{E}}|\hat{d}_{ij}-d_{ij}|$ (26)
where $M$ is the total number of chemical bonds, $\hat{d}_{ij}$ and $d_{ij}$ are the predicted and ground interatomic distances.
@@ 4.10.3 MAE of lattice
This metric calculates the error in predicting the lattice matrices:
$\Delta_{\rm lattice}=\frac{1}{9}\sum|\hat{\bm{L}}-\bm{L}|$ (27)
where $\hat{\bm{L}}$ and $\bm{L}$ are the predicted and ground lattice matrices.
@@ 4.10.4 MAE of cell volume
$\Delta_{\rm volume}=\Bigl{\lvert}|\hat{\bm{l}}_{1}\cdot(\hat{\bm{l}}_{2}\times\hat{\bm{l}}_{3})|-|\vec{\bm{l}}_{1}\cdot(\vec{\bm{l}}_{2}\times\vec{\bm{l}}_{3})|\Bigr{\rvert}$ (28)
@@ 4.10.5 Match rate
We utilize the “Structure_matcher” function from the Pymatgen package ong2013python to compare the predicted structure with the DFT-relaxed structure. Default parameters are used for this function (ltol=0.2, stol=0.3) to ensure consistent and objective comparisons.
Data availability
The dataset for X-Mn-O is available at https://zenodo.org/records/8081655 (ref. sungwon_kim_2023_8081655 ). The dataset for Materials Project is available at https://figshare.com/articles/dataset/MPF_2021_2_8/19470599 (ref. Chen2022_19470599 ). The dataset for C2DB is available at https://cmr.fysik.dtu.dk/c2db/c2db.html. The dataset for MoS2 structures with defects is available at https://research.constructor.tech/p/2d-defects-prediction. The layered vdW crystals dataset is in-house collected and currently unpublished; access can be obtained by contacting Dr. Lei Shen with reasonable requests. Source data and a Python script to reproduce the figures in this paper are provided.
Code availability